# Efficient Large Language Models: A Survey

### v2

@@ Abstract.
Large Language Models (LLMs) have demonstrated remarkable capabilities in important tasks such as natural language understanding, language generation, and complex reasoning and have the potential to make a substantial impact on our society. Such capabilities, however, come with the considerable resources they demand, highlighting the strong need to develop effective techniques for addressing their efficiency challenges. In this survey, we provide a systematic and comprehensive review of efficient LLMs research. We organize the literature in a taxonomy consisting of three main categories, covering distinct yet interconnected efficient LLMs topics from model-centric, data-centric, and framework-centric perspective, respectively. We have also created a GitHub repository where we compile the papers featured in this survey at https://github.com/AIoT-MLSys-Lab/Efficient-LLMs-Survey, and will actively maintain this repository and incorporate new research as it emerges. We hope our survey can serve as a valuable resource to help researchers and practitioners gain a systematic understanding of the research developments in efficient LLMs and inspire them to contribute to this important and exciting field.
1. Introduction
@@
Large Language Models (LLMs) are a type of advanced AI models designed to understand and generate human languages. Recently, we have witnessed a surge in LLMs include those developed by Open AI (GPT-3 (Brown et al., 2020) and GPT-4 (OpenAI, 2023)), Google (Gemini (Team and Google, 2023), GLaM (Du et al., 2022), PaLM (Chowdhery et al., 2022), PaLM-2 (Anil et al., 2023)), Meta (LLaMA-1 (Touvron et al., 2023a) and LLaMA-2 (Touvron et al., 2023b)), and other models such as BLOOM (Scao et al., 2022), PanGu- (Ren et al., 2023b), and GLM (Zeng et al., 2022). These models have demonstrated remarkable performance across a variety of tasks such as natural language understanding (NLU), language generation, complex reasoning (Yang et al., 2023b), and domain-specific tasks related to biomedicine (He et al., 2023; Wan et al., 2023, 2022), law (Eliot, 2021) and code generation (Wei et al., 2022b; Chen et al., 2021b). Such performance breakthroughs can be attributed to their massive scales in model sizes and volumes of training data, as they contain billions or even trillions of parameters while being trained on a gigantic amount of data from diverse sources.
@@
Although LLMs are leading the next wave of AI revolution, the remarkable capabilities of LLMs come at the cost of their substantial resource demands (OpenAI, 2023; Du et al., 2022; Chowdhery et al., 2022; Ren et al., 2023b). Figure 1 illustrates the relationship between model performance and model training time in terms of GPU hours for LLaMA series, where the size of each circle is proportional to the number of model parameters. As shown, although larger models are able to achieve better performance, the amounts of GPU hours used for training them grow exponentially as model sizes scale up. In addition to training, inference also contributes quite significantly to the operational cost of LLMs. Figure 2 depicts the relationship between model performance and inference throughput. Similarly, scaling up the model size enables better performance but comes at the cost of lower inference throughput (higher inference latency), presenting challenges for these models in expanding their reach to a broader customer base and diverse applications in a cost-effective way.
@@
The high resource demands of LLMs highlight the strong need to develop techniques to enhance the efficiency of LLMs. As shown in Figure 2, compared to LLaMA-1-33B, Mistral-7B (Jiang et al., 2023b), which uses grouped-query attention and sliding window attention to speed up inference, achieves comparable performance and much higher throughput. This superiority highlights the feasibility and significance of designing efficiency techniques for LLMs.
@@
where level=3text width=14em,font=,,
where level=4text width=16em,font=,,
where level=5text width=17em,font=,,
@@
Mixed Precision Acceleration, fill=blue!10
@@
Optimization Strategies, fill=blue!10
@@
Mosec,
LLM Foundry, text width=31em, fill=green!10
The overarching goal of this survey is to provide a holistic view of the technological advances in efficient LLMs and summarize the existing research directions. As illustrated in Figure 3, we organize the literature in a taxonomy consisting of three main categories, covering efficient LLMs topics from model-centric, data-centric, and framework-centric perspective, respectively. These three categories cover distinct yet interconnected research topics, collectively providing a systematic and comprehensive review of efficient LLMs research. Specifically,
Model-Centric Methods: Model-centric methods focus on both algorithm-level and system-level efficient techniques where the model itself is the focal point. With billions or even trillions of parameters, LLMs exhibit distinct characteristics (Wei et al., 2022a) compared to smaller-scale models, necessitating the development of new techniques. In §2, we survey efficient techniques that cover research directions related to model compression, efficient pre-training, efficient fine-tuning, efficient inference, and efficient architecture design.
@@
LLM Frameworks: The advent of LLMs has necessitated the development of specialized frameworks to efficiently handle their training, inference, and serving. While mainstream AI frameworks such as TensorFlow, PyTorch, and JAX provide the foundations, they lack built-in support for specific optimizations and features crucial for LLMs. In §4, we survey existing frameworks specifically designed for efficient LLMs, addressing their unique features, underlying libraries, and specializations.
In addition to the survey, we have established a GitHub repository where we compile the papers featured in the survey, organizing them with the same taxonomy: https://github.com/AIoT-MLSys-Lab/Efficient-LLMs-Survey. We will actively maintain it and incorporate new research as it emerges.
Although there are a few surveys on LLMs (Zhao et al., 2023c; Chang et al., 2023; Wang et al., 2023h; Kaddour et al., 2023), this survey provides a focused review and discussion on the literature related to the efficiency aspect of LLMs. There are also surveys on efficient Transformers (Tay et al., 2022) and their training methods (Zhuang et al., 2023a). In contrast, this survey specifically focuses on efficiency techniques designed for models of more than billions of parameters. We hope this survey together with the GitHub repo can help researchers and practitioners navigate through the literature and serve as a catalyst for inspiring further research on efficient LLMs.
2. Model-Centric Methods
2.1. Model Compression
As summarized in Figure 4, model compression techniques for LLMs can be grouped into four categories: quantization, parameter pruning, low-rank approximation, and knowledge distillation.
@@ 2.1. Model Compression
where level=3text width=17em,font=,,
@@
OBQ (Frantar and
Alistarh, 2022),
QuIP (Chee
et al., 2023),
AWQ (Lin
et al., 2023), OWQ (Lee
et al., 2023),
SpQR (Dettmers
et al., 2023b),
FineQuant (Kim
et al., 2023a), leaf, text width=18em
@@
ZeroQuant (Yao
et al., 2022a), ZeroQuant-V2 (Yao
et al., 2023b),
ZeroQuant-FP (Wu et al., 2023c),
SmoothQuant (Xiao
et al., 2023a),
@@
RPTQ (Yuan et al., 2023),
QLLM (Liu
et al., 2023b),
Outlier Suppression+ (Wei
et al., 2023),
Ahmadian et al. (2023), leaf, text width=20em
@@
LLM-QAT (Liu et al., 2023e), BitNet (Wang et al., 2023c), leaf, text width=20em
@@
LLM-Pruner (Ma et al., 2023), Sheared LLaMA (Xia
et al., 2023), LoRAPrune (Zhang et al., 2023c),leaf, text width=26em
@@
SparseGPT (Frantar and
Alistarh, 2023),
Wanda (Sun
et al., 2023),
Shao
et al. (2023), leaf, text width=21em
@@
TensorGPT (Xu et al., 2023c), LoSparse (Li
et al., 2023g), leaf, text width=15em
@@
Baby LLaMA (Timiryasov and
Tastet, 2023),
MiniLLM (Gu
et al., 2023), GKD (Agarwal et al., 2023),
KPTD (Padmanabhan et al., 2023), TED (Liang
et al., 2023), TSLD (Kim et al., 2023c),
MiniMA (Zhang
et al., 2023e), leaf, text width=23em
@@
MetaICL (Min
et al., 2021), Multitask-ICT (Huang
et al., 2022), LI et al. (2022), Lion (Jiang
et al., 2023a), DISCO (Chen et al., 2023b),
Fu
et al. (2023c),
@@
SOCRATIC CoT (Shridhar
et al., 2022),
SCOTT (Wang
et al., 2023f),
SCoTD (Li
et al., 2023a),
Peng
et al. (2023b),leaf, text width=38em
2.1.1. Quantization
@@
Quantization compresses LLMs by converting model weights and/or activations of high-precision data types such as 32-bit floating point into low-precision data types such as 8-bit integer (Dettmers et al., 2022) or 4-bit integer (Dettmers et al., 2023a):
where denotes mapping a floating number into an approximate integer; denotes the absolute maximum of the input elements; and denotes the quantization constant.
Quantization techniques for LLMs can be classified as post-training quantization (PTQ) and quantization-aware training (QAT).
@@
Post-Training Quantization (PTQ). PTQ quantizes LLMs after the model has been trained. To compensate for the accuracy drop, PTQ uses a small calibration dataset to update the quantized weights and/or activations. PTQ for LLMs can in general be grouped into two categories: weight-only quantization, and weight-activation co-quantization.
@@
Weight-Only Quantization focuses on quantizing model weights only for LLMs. For example, Dettmers et al. (2022) introduces the first multibillion-scale Int8 weight quantization method named LLM.int8 () that significantly reduces memory usage during inference while being able to maintain the full precision model performance. Frantar et al. (2023) push one step further and propose GPTQ, a post-training weight quantization method that compresses LLM weights to 3 or 4 bits instead of 8 bits. GPTQ employs layer-wise quantization with Optimal Brain Quantization (OBQ) (Frantar and Alistarh, 2022) to update weights with inverse Hessian information. This technique enables quantizing GPT models with 175 billion parameters in roughly four GPU hours with minimal accuracy loss compared to the original model. Driven by the insights that quantization can be more effective when model weights and proxy Hessian matrices are incoherent, Chee et al. (2023) propose QuIP, a post-training quantization method that applies incoherence processing to quantize LLMs to 2 bits per weight. Lin et al. (2023) observe that there exists a small portion of model weights with larger activation magnitudes referred to as salient weights that determine the quantization loss. Based on this observation, they propose a weight quantization approach named activation-aware weight quantization (AWQ) to quantize LLMs while preserving the salient weights in high precision. Similarly, Lee et al. (2023) also observe that activation outliers amplifies weight quantization loss. They propose outlier-aware weight quantization (OWQ) to identify those vulnerable weights with activation outliers and allocate high-precision to them. Dettmers et al. (2023b) propose Sparse-Quantized Representation (SpQR) to separate outlier weights that are prone to large quantization errors. These outlier weights are stored at higher precision levels, while the rest are compressed to 3-4 bits. They then propose a decoding scheme designed for the SpQR format, which accelerates the inference process on a token-by-token basis. Kim et al. (2023a) tackle the problem of outliers skewing the distribution of quantized weights, and propose FineQuant which employs an empirically crafted, heuristic-based approach to allocate varying levels of granularity to different weight matrices within the model.
@@
Weight-Activation Co-Quantization quantizes both model weights and activations. Due to the existence of outliers, activations are more difficult to quantize than model weights (Bondarenko et al., 2021). Yao et al. (2022a) propose ZeroQuant, which utilizes group-wise quantization for model weights and token-wise quantization for activations. However, ZeroQuant could not maintain accuracy for models with more than 175 billion parameters. To address this issue, Yao et al. (2023b) and Wu et al. (2023c) propose ZeroQuant-FP and ZeroQuant-V2 respectively which both utilize low-rank matrices to recover the accuracy drop. Xiao et al. (2023a) propose SmoothQuant which introduces a per-channel scaling transformation that migrates the quantization difficulty from activations to weights to achieve lossless quantization of weights and activations to 8 bits for LLMs up to 530 billion parameters. Guo et al. (2023) pinpoint outliers are critical in weight and activation quantization but their nearby normal values are not. Based on this observation, they propose OliVe, which prunes normal values adjacent to the outliers so that the outliers can be encoded with low precision. Yuan et al. (2023) identify the challenge of quantizing activations when different channels have disparate ranges. They propose RPTQ, which groups channels in activations that display similar value ranges and applies uniform quantization parameters to the values in each group. Liu et al. (2023b) propose QLLM, an adaptive channel reassembly method that efficiently tackles activation outliers and utilizes calibration data to offset the information loss incurred from quantization. Wei et al. (2023) observe that the activation outliers in LLMs are asymmetric and tend to cluster in particular channels. Based on this observation, they propose Outlier Suppression+, which introduces operations that shift and scale channels individually to neutralize asymmetric outliers. Lastly, Ahmadian et al. (2023) demonstrate that it is possible to suppress large activation outliers at scales as large as 52B. Given the right optimization choices during pre-training, they can quantize models ranging in size from 410M to 52B with minimal accuracy degradation.
@@
Quantization-Aware Training (QAT). QAT quantizes LLMs during the training process itself so as to allow LLMs to learn quantization-friendly representations. Compared to PTQ, since QAT requires training using the complete training set to make up for its accuracy drop, it is much more expensive and time consuming. Tao et al. (2022) aim to address quantization challenges in models like GPT-2 caused by uniform word embeddings, and propose QuantGPT, which combines contrastive distillation from a full-precision teacher model and logit distillation to a quantized student model during auto-regressive pretraining. LLM-QAT (Liu et al., 2023e) uses data generated by LLMs itself to distill knowledge, with the aim of quantizing a student model. Specifically, it retains the original output distribution and is capable of quantizing any generative model, irrespective of its initial training data. Besides quantizing weights and activations, LLM-QAT also tackles the quantization of the key-value cache, a crucial step for enhancing throughput and accommodating long sequence dependencies in LLMs. BitNet (Wang et al., 2023c) pioneers QAT for 1-bit LLMs, using low-precision binary weights and quantized activations, while keeping optimizer states and gradients high-precision during training, requiring only a replacement of the nn.Linear layer to train 1-bit weights from scratch.
2.1.2. Parameter Pruning

### v4

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable capabilities in important tasks such as natural language understanding and language generation, and thus have the potential to make a substantial impact on our society. Such capabilities, however, come with the considerable resources they demand, highlighting the strong need to develop effective techniques for addressing their efficiency challenges. In this survey, we provide a systematic and comprehensive review of efficient LLMs research. We organize the literature in a taxonomy consisting of three main categories, covering distinct yet interconnected efficient LLMs topics from model-centric, data-centric, and framework-centric perspective, respectively. We have also created a GitHub repository where we organize the papers featured in this survey at https://github.com/AIoT-MLSys-Lab/Efficient-LLMs-Survey. We will actively maintain the repository and incorporate new research as it emerges. We hope our survey can serve as a valuable resource to help researchers and practitioners gain a systematic understanding of efficient LLMs research and inspire them to contribute to this important and exciting field.
1 Introduction
@@
Large Language Models (LLMs) are a type of advanced AI models designed to understand and generate human languages. Recently, we have witnessed a surge in LLMs including those developed by Open AI (GPT-4 (Achiam et al., 2023) and GPT-3 (Brown et al., 2020)), Meta (LLaMA-3 (Meta, 2024), LLaMA-2 (Touvron et al., 2023b), LLaMA-1 (Touvron et al., 2023a)), and Google (Gemini (Team & Google, 2023), PaLM-2 (Anil et al., 2023), PaLM (Chowdhery et al., 2022), GLaM (Du et al., 2022)) as well as many other models such as BLOOM (Scao et al., 2023), PanGu- (Ren et al., 2023b), and GLM (Zeng et al., 2023). These models have demonstrated remarkable performance across a variety of tasks such as natural language understanding (NLU), language generation, complex reasoning (Yang et al., 2024), and domain-specific tasks related to biomedicine (He et al., 2023; Wan et al., 2023; 2022), law (Eliot, 2021) and code generation (Wei et al., 2022b; Chen et al., 2021b). Such performance breakthroughs can be attributed to their massive scales in model sizes and volumes of training data, as they contain billions or even trillions of parameters while being trained on a gigantic amount of data from diverse sources.
@@
Although LLMs are leading the next wave of AI revolution, their remarkable capabilities come at substantial resource demands (Achiam et al., 2023; Du et al., 2022; Chowdhery et al., 2022; Ren et al., 2023b). Figure 1 illustrates the relationship between model performance and model training time in terms of GPU hours for LLaMA series, where the size of each circle is proportional to the number of model parameters. As shown, although larger models are able to achieve better performance, the amounts of GPU hours used for training them grow exponentially as model sizes scale up. In addition to training, inference also contributes quite significantly to the operational cost of LLMs. Figure 2 depicts the relationship between model performance and inference throughput. Similarly, scaling up the model size enables better performance but comes at the cost of lower inference throughput (higher inference latency), presenting challenges for these models in expanding their reach to a broader customer base and diverse applications in a cost-effective way.
@@
The high resource demands of LLMs highlight the strong need to develop techniques to enhance the efficiency of LLMs. As shown in Figure 2, compared to LLaMA-1-33B, Mistral-7B (Jiang et al., 2023a), which uses grouped-query attention and sliding window attention to speed up inference, achieves comparable performance and much higher throughput. This superiority highlights the feasibility and significance of designing efficiency techniques for LLMs.
@@
where level=3text width=16em,font=,,
where level=4text width=18em,font=,,
where level=5text width=18em,font=,,
@@
Mixed Precision Training, fill=blue!10
@@
Training Optimizers, fill=blue!10
@@
Mosec, text width=43em, fill=green!10
The overarching goal of this survey is to provide a holistic view of the technological advances in efficient LLMs. As illustrated in Figure 3, we organize the literature in a taxonomy consisting of three main categories, covering efficient LLMs topics from model-centric, data-centric, and framework-centric perspective, respectively. These three categories cover distinct yet interconnected research topics, collectively providing a systematic and comprehensive review of efficient LLMs research. Specifically,
Model-Centric Methods: Model-centric methods focus on both algorithm-level and system-level efficient techniques where the model itself is the focal point. With billions or even trillions of parameters, LLMs exhibit distinct characteristics (Wei et al., 2022a) compared to smaller-scale models, necessitating the development of new techniques to enhance their efficiency. In §2, we survey efficient techniques that cover research directions related to model compression, efficient pre-training, efficient fine-tuning, efficient inference, and efficient architecture design.
@@
LLM Frameworks: The advent of LLMs necessitates the development of specialized frameworks to efficiently handle their training, fine-tuning, inference, and serving. While mainstream AI frameworks such as TensorFlow and PyTorch provide the foundations, they lack built-in support for specific optimizations and features crucial for LLMs. In §4, we survey existing frameworks specifically designed for efficient LLMs, covering their unique features, underlying libraries, and specializations.
In addition to the survey, we have established a GitHub repository where we compile the papers featured in this survey at https://github.com/AIoT-MLSys-Lab/Efficient-LLMs-Survey. We will actively maintain it and incorporate new research as it emerges.
Although there are a few surveys on LLMs (Zhao et al., 2023a; Chang et al., 2024; Wang et al., 2023i; Kaddour et al., 2023), this survey provides a focused review and discussion on the literature related to the efficiency aspect of LLMs. There are also surveys on efficient Transformers (Tay et al., 2022) and their training methods (Zhuang et al., 2023). In contrast, this survey specifically focuses on efficiency techniques designed for models of more than billions of parameters. We hope this survey together with the GitHub repository can help researchers and practitioners navigate through the literature and serve as a catalyst for inspiring further research on efficient LLMs.
2 Model-Centric Methods
2.1 Model Compression
Model compression enhances efficiency by reducing the sizes and the amount of arithmetic operations of LLMs. Unlike conventional model compression techniques, most LLM compression approaches are designed under the post-training setting to avoid resource-intensive retraining. As summarized in Figure 4, model compression techniques for LLMs can be grouped into four categories: quantization, parameter pruning, low-rank approximation, and knowledge distillation. These four categories are orthogonal to each other, and compress LLMs from different perspectives.
@@ 2.1 Model Compression
where level=3text width=18em,font=,,
@@
OBQ (Frantar & Alistarh, 2022),
QuIP (Chee et al., 2023),
AWQ (Lin et al., 2023), OWQ (Lee et al., 2023),
SpQR (Dettmers et al., 2024),
FineQuant (Kim et al., 2023d), leaf, text width=30em
@@
ZeroQuant (Yao et al., 2022b),
ZeroQuant-FP (Wu et al., 2023b),
ZeroQuant-V2 (Yao et al., 2023d),
SmoothQuant (Xiao et al., 2023),
@@
RPTQ (Yuan et al., 2023a),
Ahmadian et al. (2023),
Outlier Suppression+ (Wei et al., 2023),
QLLM (Liu et al., 2024c), leaf, text width=31em
@@
LLM-QAT (Liu et al., 2023d), BitNet (Wang et al., 2023b), leaf, text width=40em
@@
LLM-Pruner (Ma et al., 2023), Sheared LLaMA (Xia et al., 2023), LoRAPrune (Zhang et al., 2023c),
LoRAShear (Chen et al., 2023e), Deja Vu (Liu et al., 2023f),leaf, text width=45em
@@
SparseGPT (Frantar & Alistarh, 2023),
Wanda (Sun et al., 2024),
Shao et al. (2024), leaf, text width=38em
@@
TensorGPT (Xu et al., 2023a), LoSparse (Li et al., 2023g), FWSVD (Hsu et al., 2022), ASVD (Yuan et al., 2023b), SVD-LLM (Wang et al., 2024c), leaf, text width=65.5em
@@
Baby LLaMA (Timiryasov & Tastet, 2023),
MiniLLM (Gu et al., 2024),
KPTD (Padmanabhan et al., 2023),
TED (Liang et al., 2023), TSLD (Kim et al., 2023b),
MiniMA (Zhang et al., 2023a),
GKD (Agarwal et al., 2024), leaf, text width=50em
@@
MetaICL (Min et al., 2022a),
Multitask-ICT (Huang et al., 2022),
Li et al. (2024b),
Lion (Jiang et al., 2023b),
DISCO (Chen et al., 2023j),
Fu et al. (2023b),
@@
SOCRATIC CoT (Shridhar et al., 2023),
SCOTT (Wang et al., 2023c),
SCoTD (Li et al., 2023b),
Peng et al. (2023a), Zephyr (Tunstall et al., 2023),leaf, text width=49em
2.1.1 Quantization
@@
Quantization compresses LLMs by converting model weights and/or activations of high-precision data types such as 32-bit floating point into low-precision data types such as 8-bit integer (Dettmers et al., 2023) as:
where denotes mapping a floating point number into an approximate integer; denotes the absolute maximum of the input elements; and denotes the quantization constant. Quantization techniques for LLMs can be classified into post-training quantization (PTQ) and quantization-aware training (QAT). Compared to other LLM compression methods such as parameter pruning and low-rank approximation, quantization methods have been shown to achieve superior compression-accuracy trade-offs (Li et al., 2024c).
@@
Post-Training Quantization (PTQ). PTQ quantizes LLMs after the model has been trained. To compensate for the accuracy drop, PTQ uses a small calibration dataset to update the quantized weights and/or activations. PTQ in general can be grouped into two categories: weight-only quantization, and weight-activation co-quantization.
@@
Weight-Only Quantization focuses on quantizing model weights only. For example, Dettmers et al. (2022) introduce the first multi-billion-scale 8-bit integers (or INT8) weight quantization method named LLM.int8() that significantly reduces memory usage during inference while being able to maintain the performance of the full-precision model. Frantar et al. (2023) push one step further and propose GPTQ, a post-training weight quantization method that compresses LLM weights to 3 or 4 bits instead of 8 bits. GPTQ employs layer-wise quantization with Optimal Brain Quantization (OBQ) (Frantar & Alistarh, 2022) to update weights with inverse Hessian information. This technique enables quantizing GPT models with 175 billion parameters in roughly four GPU hours with minimal accuracy drop compared to the original model. Furthermore, driven by the insights that quantization can be more effective when model weights and proxy Hessian matrices are incoherent, Chee et al. (2023) propose QuIP, a post-training quantization method that applies incoherence processing to quantize LLMs to 2 bits per weight. As another line of research under weight-only quantization, Lin et al. (2023) observe that there exists a small subset of model weights, characterized by larger activation magnitudes, known as salient weights, play a crucial role in determining the quantization loss. Based on this observation, they propose an approach named activation-aware weight quantization (AWQ) to quantize LLMs while preserving the salient weights in high precision, demonstrating superior performance over GPTQ. Similarly, Lee et al. (2023) observe that activation outliers amplify weight quantization loss. They propose outlier-aware weight quantization (OWQ) to identify those vulnerable weights with activation outliers and allocate high-precision to them. Dettmers et al. (2024) introduce Sparse-Quantized Representation (SpQR) to separate outlier weights that are prone to large quantization errors. These outlier weights are preserved at higher precision, while the remaining weights are compressed to 3-4 bits. Additionally, they introduce a decoding scheme tailored for the SpQR format that enhances the efficiency of inference on a token-by-token basis. Lastly, Kim et al. (2023d) aim to address the issue of outliers that distort the distribution of quantized weights, and propose FineQuant that employs an empirically crafted, heuristic-based approach to allocate varying levels of granularity to different weight matrices.
Figure 5: Illustrations of model compression techniques for LLMs. -
@@
Weight-Activation Co-Quantization differs from weight-only quantization in the sense that it quantizes both model weights and activations. For instance, Yao et al. (2022b) propose ZeroQuant, which combines group-wise quantization for model weights and token-wise quantization for activations. However, ZeroQuant falls short in maintaining accuracy for models with more than 175 billion parameters. To address this issue, Yao et al. (2023d) and Wu et al. (2023b) propose ZeroQuant-FP and ZeroQuant-V2 respectively, both of which utilize low-rank matrices to recover the accuracy drop. A key challenge of weight-activation co-quantization is that due to the existence of outliers, activations are more difficult to quantize than model weights (Bondarenko et al., 2021). To address this challenge, Xiao et al. (2023) propose SmoothQuant which introduces a per-channel scaling transformation that migrates the quantization difficulty from activations to weights to achieve lossless quantization of weights and activations to 8 bits for LLMs up to 530 billion parameters. Guo et al. (2023) pinpoint that outliers are critical in weight-activation co-quantization but their nearby normal values are not. Given that, they propose OliVe, which prunes normal values adjacent to the outliers so that the outliers can be encoded with higher precision. Yuan et al. (2023a) identify the challenge of quantizing activations when different channels have disparate ranges. They propose RPTQ, which groups channels in activations that have similar value ranges and applies uniform quantization parameters to the values in each group. Ahmadian et al. (2023) demonstrate that it is possible to suppress large activation outliers at scales as large as 52B. Given the right optimization choices during pre-training, they can quantize models ranging in size from 410M to 52B with minimal accuracy degradation. Wei et al. (2023) observe that the activation outliers in LLMs are asymmetric and tend to cluster in particular channels. Based on this observation, they propose Outlier Suppression+, which introduces operations that shift and scale channels individually to neutralize asymmetric outliers. Lastly, Liu et al. (2024c) propose QLLM, an adaptive channel reassembly method that tackles activation outliers and utilizes calibration data to offset the information loss incurred from quantization. Experimental result shows that QLLM achieves better compression performance than SmoothQuant and Outlier Suppression+ on LLaMA model family.
@@
Quantization-Aware Training (QAT). Different from PTQ, QAT quantizes LLMs during the training process, allowing LLMs to learn quantization-friendly representations. Since QAT requires training using the complete training dataset, it is much more expensive and time consuming than PTQ. Tao et al. (2022) propose QuantGPT, which combines contrastive distillation from a full-precision teacher model and logit distillation to a quantized student model during autoregressive pretraining. QuantGPT achieves 14.4× and 13.4× compression rates on GPT-2 and BART with comparable performance with the full-precision models. LLM-QAT (Liu et al., 2023d) uses data generated by LLMs itself to distill knowledge with the objective of quantizing a student model. Specifically, LLM-QAT retains the original output distribution and is capable of quantizing a model irrespective of its initial training data. Besides quantizing weights and activations, LLM-QAT also quantizes the key-value cache, a crucial step for enhancing throughput and accommodating long sequence dependencies in LLMs. Experimental results show that LLM-QAT achieves better performance over training-free methods especially in low-bit settings. Lastly, BitNet (Wang et al., 2023b) pioneers QAT for 1-bit LLMs. It proposes to use low-precision binary weights and quantized activations while keeping optimizer states and gradients high-precision during training. Experimental results show that compared to FP16 Transformer baselines, BitNet is able to achieve competitive performance while substantially reducing memory footprint and energy consumption.
2.1.2 Parameter Pruning